### v1

@@ Abstract
Recent state-of-the-art object segmentation mechanisms, such as the Segment Anything Model (SAM) and FastSAM, first encode the full image over several layers and then focus on generating the mask for one particular object or area. We present an off-grid Fovea-Like Input Patching (FLIP) approach, which selects image input and encodes it from the beginning in an object-focused manner. While doing so, it separates locational encoding from an object-centric perceptual code. FLIP is more data-efficient and yields improved segmentation performance when masking relatively small objects in high-resolution visual scenes. On standard benchmarks such as Hypersim, KITTI-360, and OpenImages, FLIP achieves Intersection over Union (IoU) scores that approach the performance of SAM with much less compute effort. It surpasses FastSAM in all IoU measurements. We also introduce an additional semi-natural but highly intuitive dataset where FLIP outperforms SAM and FastSAM overall and particularly on relatively small objects. Seeing that FLIP is an end-to-end object-centric segmentation approach, it has high potential particularly for applications that benefit from computationally efficient, spatially highly selective object tracking.
@@ 1 Introduction
Figure 1: Comparison of segmentation performance (Mean IoU) on our ObjaScale dataset versus model size (number of parameters, in millions). FLIP variants achieve superior performance with significantly fewer parameters compared to SAM variants.
@@
The FastSAM model Zhao et al. (2023) addresses the former limit. FastSAM replaces the first stage transformer-based encoder architecture in SAM with a convolutional ANN approach (CNN), which is pre-trained to segment the whole image. The query then targets the latent encodings and produces one area- or object-specific mask selectively. As a result, FastSAM yields performance close-to SAM, but is trained on only 2% of the original SAM dataset, has much fewer parameters, and partially runs an order of magnitude faster.
The latter limit remains an open challenge. At this point nearly all segmentation techniques, including SAM and FastSAM, rely on a full image encoder. This is also the case for most object centric models, such as SlotAttention, VideoSAUR, and related work Locatello et al. (2020); Elsayed et al. (2022); Singh et al. (2022); Zadaianchuk et al. (2024), which first encode the entire image before assigning information to slots. The challenge to computationally efficiently segment and track small but potentially high-resolution objects across diverse and complex scenes remains.
To this end, we introduce FLIP: a fovea-like input patching approach that is integrated in an object-centric, off-grid vision framework. FLIP dynamically adapts its processing pipeline to the object’s size and spatial characteristics. It ignores currently irrelevant image subregions and focusses on critical regions with a flexible, multi-resolution approach. Our key contributions are:
  • Off-Grid, Scale-Invariant Object Encoding: We introduce a fovea-inspired patch sampling method that directly encodes image regions off-grid, adaptively focusing on objects of interest in a multi-resolution fashion. This scale-invariant approach is robust to large variations in object size and image resolution. It enables the detailed encoding of very small objects even in high-resolution scenes.
  • Interpretable Object-Centric Latent Representation: Building on insights from cognitive science, FLIP explicitly segregates perceptual features from positional cues into distinct latent embeddings. This structured representation not only boosts segmentation accuracy but also improves explainability, as it allows for a separated assessment of object identity and spatial placement.
  • State-of-the-Art Segmentation Performance with High Parameter Efficiency: Despite using significantly fewer parameters compared to state-of-the-art models like SAM or FastSAM, FLIP achieves competitive (SAM) or superior (FastSAM) segmentation accuracy on benchmarks such as Hypersim, KITTI-360, and OpenImages.
  • Superior performance on novel dataset ObjaScale For further evaluation, we present a segmentation dataset that contains high-resolution real-world backgrounds combined with objects of varying scales, rendered in Blender. FLIP outperforms all other methods on ObjaScale, even with the smallest parameter settings (Figure 1). Note that ObjaScale is used for evaluation only, neither FLIP nor SAM or FastSAM were trained on it.
Figure 2: The multi-resolution patch sampling procedure in FLIP sequentially samples tiny (blue), small (green), and large (red) patches given an object-centered 2D Gaussian. Overlapping patches are prohibited, thus generating a fovea-like distribution of patches across multiple resolutions.
@@ 3 Methods
Figure 4: Overview of the FLIP architecture. The Foveal Patching module dynamically samples multi-resolution patches centered around objects of interest. These patches are embedded into a unified latent space using resolution-specific Patch Embedding Modules ($E_{r_{0}})$ to $E_{r_{K}}$). The Vision Transformer Encoder processes the embedded patches, generating separate Perceptual ($z_{p}$) and Positional Codes ($z_{\text{pos}}$). The ViT predictor combines these codes to predict the object instance segmentation mask, reconstructing it via Patch De-embedding Modules ($D_{r_{0}}$ to $D_{r_{K}})$).
In this section, we present the overall Fovea-Like Input Patching (FLIP) architecture. Its main processing pipeline is shown in Figure 4. FLIP is a supervised vision model for efficient object-centric segmentation. It combines a fovea-inspired patching mechanism with a Vision Transformer (ViT)-based encoder. The encoded information is then used to generate the targeted object’s segmentation mask, location, and bounding box, thereby disentangling perceptual and positional information.
The fovea-like patching mechanism dynamically selects multi-resolution patches based on a 2D Gaussian distribution, which has an effect similar to the query in SAM and FastSAM but acts directly on the input image. High-resolution patches focus on the object center, capturing fine details, while coarser patches cover peripheral regions, which inform FLIP about the surrounding context.
The ViT encoder processes the sampled patches and outputs two representations: a perceptual code ($z_{p}$), which captures the object’s appearance and shape, and a positional code ($z_{\text{pos}}$), which encodes location, orientation, and scale.
The predictor combines $z_{p}$ and $z_{\text{pos}}$ to predict the segmentation mask. The positional code guides dynamic resolution-dependent decoding, ensuring high-resolution predictions at the object’s spatial extent, while the perceptual code provides detailed object-specific information. This design enables FLIP to excel at segmenting small objects in high-resolution scenes with much lower computational overhead when compared to SAM and FastSAM.
FLIP is trained end-to-end with supervised losses for mask prediction, bounding box localization, and positional accuracy. In the following subsections, we provide the mathematical details of the fovea-like sampling mechanism, the ViT encoder, and the ViT predictor.
@@ 3.1 Fovea-like Input Patching
We equip our ViT-based model with a multi-resolution, fovea-inspired patching mechanism that centers around the object of interest, thereby preserving high-resolution detail at the object center and coarser coverage in peripheral regions. Specifically, we derive the object center $\mu=(\mu_{x},\mu_{y})$ and covariance $\Sigma$ from the ground-truth mask, yielding a 2D Gaussian $\mathcal{N}(\mu,\Sigma)$ that approximates the object’s spatial extent and orientation in compressed form. From $\Sigma$, we extract rotation $(\theta_{a},\theta_{b})$ and scale $(\sigma_{x},\sigma_{y})$ via a standard eigenvalue decomposition of the covariance matrix. For input sampling, the Gaussian serves as a spatial input query—similar to the prompts in SAM and FastSAM—from which we then draw patches at multiple scales.
@@
We define $K$ patch sizes $p_{1}<p_{2}<\dots<p_{K}$, from smallest (highest resolution) to largest. We fix the total number of patches $N$ and choose $N_{i}$ patches for patch size $p_{i}$ as follows. First, we compute default numbers $\hat{N}_{i}$ by approximating the area covered by the Gaussian’s 99% inner density mass from $\Sigma$ and dividing it by $p_{i}^{2}$. Then, starting at the coarsest resolution (i.e., $p_{K}$), we compute $N_{i}=min(\hat{N}_{i},N-\sum_{j=i+1}^{K}N_{j})$. As a result, all chosen patch sizes are smaller than the Gaussian inner area and the number of chosen patches per size distributes itself from coarse to fine until the total number of patches $N$ is reached($N=\sum_{i}^{K}N_{i}$).
@@
We sample $N_{i}$ patches from $\mathcal{N}(\mu,\Sigma)$. Higher-resolution patches (small $p_{i}$) are drawn first. Next, progressively larger patches (lower resolution) are sampled from the remaining density excluding the patches that overlap with patches that were already chosen (cf. Figure 2). Each sampled patch is first flattened and then mapped to a common embedding space via resolution-specific encoders $E_{r_{i}}$ (similarly to Jaegle et al., 2021), yielding a set of tokens $T=\{t_{1},...,t_{N}\}$ of equal tensor size. These embedded tokens are concatenated and fed to the main ViT layers. Note that because sampling depends on $\mu$ and $\Sigma$ only, the approach is in principle independent of the full image size.
3.2 Object-Centric Attention
For further encoding, we use modified Transformer blocks that preserve the “what vs. where” separation. Figure 5 shows the design of one block. Inspired by Yu et al. (2023), we inject positional embeddings in every layer, but only into the queries and keys (never the values), ensuring spatial cues influence attention without affecting perceptual content. The first layer $l=1$ block receives as input the set of tokens $T$, that is, $X_{1}=T$. Further, all blocks receive the following positional encoding $PE$, in the form of relative positional embeddings. We compute these embeddings relative to the 2D Gaussian input query, shifting each token coordinate $(x_{i},y_{i})$ by $(\mu_{x},\mu_{y})$, rotating it by $(\theta_{a},\theta_{b})$, and scaling it by $(\sigma_{x},\sigma_{y})$. We then generate the embedding $\mathrm{PE}$ via an MLP. This embedding is then added to queries and keys:
$\mathrm{PE}_{i}$ $=MLP_{i}(\hat{x}_{i},\hat{y}_{i}),\quad\mathrm{PE}_{j}=MLP_{j}(\hat{x}_{j},\hat{y}_{j}),$
$\mathbf{Q}^{\prime}_{i}$ $=\mathbf{Q}_{i}+\mathrm{PE}_{i},\quad\mathbf{K}^{\prime}_{j}=\mathbf{K}_{j}+\mathrm{PE}_{j}.$
We adopt a pre-layer normalization scheme Xiong et al. (2020), such that LayerNorm precedes both the self-attention and FFN blocks. Consequently, the Transformer layer computations for input $X_{l}$ proceed as:
$\hat{X}$ $=\mathrm{LN}(X_{l}),$
$Q$ $=\hat{X}W_{Q},\quad K=\hat{X}W_{K},\quad V=\hat{X}W_{V},$
$Q^{\prime}$ $=Q+\mathrm{PE},\quad K^{\prime}=K+\mathrm{PE},$
$Y$ $=X_{l}+\mathrm{softmax}\left(\frac{Q^{\prime}K^{\prime T}}{\sqrt{d_{k}}}\right)V,$
$X_{l+1}$ $=Y+\mathrm{FFN}(\mathrm{LN}(Y)),$
where $d_{k}$ is the key dimension. By avoiding positional biases in $V$, the model attends spatially via $\mathrm{PE}$ without affecting the perceptual content. This design aligns with our object-centric principle: positional information modulates attention weights (queries, keys), and perceptual information flows through $V$. Depending on the chosen network sizes, we use three, six, or seven of these encoding blocks.
@@ 3.2 Object-Centric Attention
Figure 5: Architecture of the Object-Centric Attention block in FLIP. Positional embeddings influence only the queries and keys, maintaining separation of spatial and perceptual information for efficient object-centric processing.
3.3 Learning What and Where
Following the Loci family of models Traub et al. (2023; 2024a; 2024b), which draw inspiration from the dorsal-ventral visual pathway separation in the brain, FLIP finally enforces a strict separation between perceptual and positional information at the end of the encoding pipeline. This separation ensures that the model independently processes “what” an object is and “where” it is located, encouraging a more interpretable object segmentation. The input to this separation comes from the output of the last object-centric attention block, which can be interpreted as consisting of $N$ encoded tokens $t_{i}$ and associated image coordinates $(x_{i},y_{i})$.
Perception Code $\mathbf{z}_{p}$.
To generate the final perception code $\mathbf{z}_{p}$, we combine all tokens by average pooling them and refining them via a cross-attention-layer, using the average pooled value as the query:
$\mathbf{z}_{avg}$ $=\tfrac{1}{N}\!\sum_{i=1}^{N}\mathbf{t}_{i}$
$\mathbf{z}_{p}$ $=\mathbf{z}_{avg}+\text{CrossAttentionLayer}(\mathbf{z}_{avg},\,T).$
This yields a compact representation of the object’s appearance.
Positional Code $\mathbf{z}_{\text{pos}}$.
To infer $\mathbf{z}_{\text{pos}}$, cross-attention modules on token coordinates obtain object-focused mean position, scale, and orientation estimates. First, we compute the object center:
$\mathbf{q}_{\mu}=W_{q}^{\mu}\,\mathbf{1},\quad\mathbf{k}_{\mu,i}=W_{k}^{\mu}\,t_{i},\quad a_{\mu,i}=\text{softmax}\bigl{(}\mathbf{q}_{\mu}\!\cdot\!\mathbf{k}_{\mu,i}^{\top}\bigr{)},$
$\mu_{x}=\sum_{i=1}^{N}a_{\mu,i}\,x_{i},\quad\mu_{y}=\sum_{i=1}^{N}a_{\mu,i}\,y_{i}.$
Similarly, additional cross-attention heads estimate $\sigma_{x},\sigma_{y}$ (spread) and $\theta_{a},\theta_{b}$ (rotation). Concatenating these yields:
$\mathbf{z}_{\text{pos}}=[\mu_{x},\,\mu_{y},\,\sigma_{x},\,\sigma_{y},\,\theta_{a},\,\theta_{b}],$
which enables us in principle to use this positional output also as an input query. This option is however not further explored in this paper.
3.4 Mask Prediction
To generate mask predictions, we first choose the region of the image that is expected to contain the object. FLIP then generates output patches that cover this region, predicting the mask via the patches. To determine the output region we take the positional code $\mathbf{z}_{\text{pos}}$ and compute the rotated coordinates via:
$x_{\text{rot}}$ $=\theta_{a}\,(x-\mu_{x})-\theta_{b}\,(y-\mu_{y})$
$y_{\text{rot}}$ $=\theta_{b}\,(x-\mu_{x})+\theta_{a}\,(y-\mu_{y}),$
with $\sqrt{\theta_{a}^{2}+\theta_{b}^{2}}=1$ and $\sigma_{x},\sigma_{y}$ clipped to $\geq\epsilon$. The Gaussian map is
$G(x,y)=\exp\!\Bigl{(}-\frac{1}{2}\Bigl{[}\bigl{(}\frac{x_{\text{rot}}}{\sigma_{x}}\bigr{)}^{2}+\bigl{(}\frac{y_{\text{rot}}}{\sigma_{y}}\bigr{)}^{2}\Bigr{]}\Bigr{)},$
thresholded at $0.01$ (its 99% confidence region).
We fully cover this region, by selecting $N_{out}$ patches using two resolutions. The patch ratio for these resolutions is chosen adaptively from $\sigma_{x},\sigma_{y}$ to guarantee that exactly $N_{out}$ patches are needed to cover the determined output region. For each patch $(x_{i},y_{i})$ we form an output token:
$t_{\mathrm{out},i}=\mathbf{z}_{p}+\mathrm{PE}(x_{i},y_{i}),$
where $\mathbf{z}_{p}$ is the perception code. These tokens are then processed through two Transformer blocks, whose block structure is identical to the encoder blocks. Finally, a resolution-specific de-embedding module $D_{r_{i}}$ reconstructs each patch $\hat{P}_{i}$:
$\hat{P}_{i}=D_{r_{i}}\bigl{(}t_{\mathrm{out},i}\bigr{)}.$
By placing each patch $\hat{P}_{i}$ at its respective position $(x_{i},y_{i})$ we cover the output region with patch-respective mask predictions and set the remaining area to zero (i.e., not part of the mask). This concludes the mask prediction process.

### v2

@@ Abstract
Current state-of-the-art segmentation models encode entire images before focusing on specific objects. As a result, they waste computational resources—particularly when small objects are to be segmented in high-resolution scenes. We introduce FLIP (Fovea-Like Input Patching), a parameter-efficient vision model that realizes object segmentation through biologically-inspired top-down attention. FLIP selectively samples multi-resolution patches centered on objects of interest from the input. As a result, it allocates high-resolution processing to object centers while maintaining coarser peripheral context. This off-grid, scale-invariant design enables FLIP to outperform META’s Segment Anything models (SAM) by large margins: With more than 1000x fewer parameters, FLIP-Tiny (0.51M parameters) reaches a mean IoU of 78.24% while SAM-H reaches 75.41% IoU (641.1M parameters). FLIP-Large even achieves 80.33% mean IoU (96.6M parameters), still running about 6$\times$ faster than SAM-H. We evaluate on six benchmarks in total. In five established benchmarks (Hypersim, KITTI-360, OpenImages, COCO, LVIS) FLIP consistently outperforms SAM and various variants of it. In our novel ObjaScale dataset, which stress-tests scale invariance with objects ranging from 0.0001% up-to 25% of the image area, we show that FLIP segments even very small objects accurately, where existing models fail severely. FLIP opens new possibilities for real-time, object-centric vision applications and offers much higher energy efficiency. We believe that FLIP can act as a powerful foundation model, as it is very well-suited to track objects over time, for example, when being integrated into slot-based scene segmentation architectures.
@@ 1 Introduction
Figure 1: Comparison of mean IoU across six diverse datasets (Hypersim, KITTI-360, OpenImages, COCO, LVIS, ObjaScale) plotted against model size (in millions of parameters). FLIP variants consistently outperform SAM variants while using several orders of magnitude fewer parameters.
@@
Derivative model variants like EfficientSAM, MobileSAM, and FastSAM Xiong et al. (2023); Zhang et al. (2023); Zhao et al. (2023) address the former limitation. In our evaluation of single-object segmentation performance across multiple datasets, EfficientSAM employs SAMI (SAM-leveraged masked image pretraining), training lightweight ViT encoders to reconstruct features from SAM’s ViT-H encoder, and achieves 72.29% mean IoU with only 10.22M parameters (EfficientSAM-T) compared to SAM-H’s 75.41% IoU with 641M parameters. MobileSAM applies decoupled distillation, replacing SAM’s ViT-H encoder (632M parameters) with TinyViT (5M parameters) Wu et al. (2022), achieving 71.33% mean IoU with an over 60$\times$ parameter reduction and $\sim$21ms runtime per image. FastSAM replaces the first-stage transformer-based encoder in SAM with a convolutional ANN (CNN) that is pre-trained to segment full images. It achieves a runtime per image of under 10ms but exhibits significantly reduced performance with only 44.58% mean IoU (FastSAM-s).
At this point nearly all segmentation techniques, including SAM and its efficient variants, rely on a full image encoder. This is also the case for most object centric models, such as SlotAttention, VideoSAUR, and related work Locatello et al. (2020); Elsayed et al. (2022); Singh et al. (2022); Zadaianchuk et al. (2024), which first encode the entire image before assigning information to slots. The challenge to computationally efficiently segment and track small but potentially high-resolution objects across diverse and complex scenes remains. Here, we introduce a vision transformer-based foundational model that utilizes a local, pixel-patch-based image encoder to generate accurate masks.
Figure 3: Overview of the FLIP architecture. The Foveal Patching module dynamically samples multi-resolution patches centered around objects of interest. These patches are embedded into a unified latent space using resolution-specific Patch Embedding Modules ($E_{r_{0}})$ to $E_{r_{K}}$). The Vision Transformer Encoder processes the embedded patches, generating keys $K_{1..n}$ and values $V_{1..n}$. The Pixel-Predictor performs attention over queries derived from pixel coordinates $Q_{x,y}$ enabling instance segmentation with pixel-level precision.
In particular, we introduce FLIP: a fovea-like input patching approach that is integrated in an object-centric, off-grid vision framework. FLIP dynamically adapts its processing pipeline to the object’s size and spatial characteristics. It ignores currently irrelevant image subregions and focuses on critical regions with a flexible, multi-resolution approach. Our key contributions are:
  • Off-Grid, Scale-Invariant Object Encoding: We introduce a fovea-inspired pixel-based patch sampling method that directly encodes image regions off-grid, adaptively focusing on objects of interest in a multi-resolution fashion. This scale-invariant approach is robust to large variations in object size and image resolution. It enables the detailed encoding of very small objects even in high-resolution scenes.
  • Exceptional Parameter and Computational Efficiency: FLIP achieves superior segmentation performance while using orders of magnitude fewer parameters than existing state-of-the-art models. With lightweight encoders and efficient processing, FLIP provides significant energy savings and faster inference times, making it suitable for real-time applications and resource-constrained environments.
  • State-of-the-Art Segmentation Performance with High Parameter Efficiency: Despite using significantly fewer parameters compared to state-of-the-art models from the SAM family, FLIP achieves superior segmentation accuracy on standard benchmarks such as Hypersim, KITTI-360, OpenImages, COCO and LVIS.
  • Superior performance on novel dataset ObjaScale We introduce ObjaScale, a high-resolution evaluation benchmark with Blender-rendered objects at varying scales on real-world High Dynamic Range Image (HDRI) backgrounds, designed to stress-test scale invariance. As in the other datasets, FLIP outperforms all SAM variants on ObjaScale and shows that it can segment even rather tiny objects accurately.
@@ 3 Methods
In this section, we present the overall Fovea-Like Input Patching (FLIP) architecture. Its main processing pipeline is shown in Figure 3. FLIP is a supervised vision model for efficient object-centric segmentation. It combines a fovea-inspired patching mechanism with a Vision Transformer (ViT)-based encoder. The encoded information is then used to generate the targeted object’s segmentation mask.
The fovea-like patching mechanism dynamically selects multi-resolution patches based on a 2D Gaussian distribution, which has an effect similar to the query in SAM variants but acts directly on the input image. High-resolution patches focus on the object center, capturing fine details, while coarser patches cover peripheral regions, which inform FLIP about the surrounding context.
The ViT encoder processes the sampled patches and outputs latent tokens from which we compute keys $K_{1..n}$ and values $V_{1..n}$. The Pixel-Predictor then computes queries $Q_{x,y}$ from pixel coordinates and uses an attention mechanism to predict the segmentation mask with pixel level accuracy (see Figure 5).
FLIP is trained end-to-end with a sparse supervised loss for mask prediction. In the following subsections, we provide the mathematical details of the fovea-like sampling mechanism, the ViT encoder, and the Pixel-Predictor.
@@ 3.1 Fovea-like Input Patching
We equip our ViT-based model with a multi-resolution, fovea-inspired patching mechanism that centers around the object of interest. The mechanism preserves high-resolution detail at the object center and coarser coverage in peripheral regions. Specifically, we derive the object center $\mu=(\mu_{x},\mu_{y})$ and covariance $\Sigma$ from the ground-truth mask, yielding a 2D Gaussian $\mathcal{N}(\mu,\Sigma)$ that approximates the object’s spatial extent and orientation in compressed form. For input sampling, the Gaussian serves as a spatial input query—similar to the prompts in SAM variants—from which we then draw patches at multiple scales.
@@
We define $K$ patch sizes $p_{1}<p_{2}<\dots<p_{K}$, from smallest (highest resolution) to largest. We fix the total number of patches $N$ and choose $N_{i}$ patches for patch size $p_{i}$ as follows. First, we compute default numbers $\hat{N}_{i}$ by approximating the integral of the 2D Gaussian and dividing by $p_{i}^{2}$, scaled by a coverage parameter $c\in(0.1,2)$ that controls the density of patch sampling. Then, starting at the coarsest resolution (i.e., $p_{K}$), we compute $N_{i}=\min(\hat{N}_{i},N-\sum_{j=i+1}^{K}N_{j})$. As a result, the chosen patches cover the Gaussian inner area with a density proportional to $c$ and the number of chosen patches per size distributes itself from coarse to fine until the total number of patches $N$ is reached ($N=\sum_{i}^{K}N_{i}$).
@@
We sample $N_{i}$ patches from $\mathcal{N}(\mu,\Sigma)$ with continuous coordinates, allowing patches to be centered at real-valued positions $(x,y)\in\mathbb{R}^{2}$. Pixel values are extracted using bilinear interpolation, enabling sub-pixel precision. To control patch overlap, we employ a spatial hash grid that tracks placed patches. When sampling, we reject candidate positions that would result in overlap exceeding the maximum overlap threshold $\tau\in[0,4]$. This ensures efficient spatial coverage while allowing for controlled redundancy. We implement this sampling mechanism as a Python C extension in the data loading pipeline, achieving approximately 1ms average extraction time per image. Each sampled patch is first flattened and then mapped to a common embedding space via resolution-specific encoders $E_{r_{i}}$ (similarly to Jaegle et al., 2021), yielding a set of tokens $T=\{t_{1},...,t_{N}\}$ of equal tensor size. These embedded tokens are concatenated and fed to the main ViT layers. Note that the approach is in principle independent of the full image size, because sampling depends on $\mu$ and $\Sigma$ only.
3.2 Encoder
For encoding the multi-resolution patches, we employ a Vision Transformer architecture with modified attention blocks inspired by LaPE Yu et al. (2023). Our design differs from standard transformers by incorporating position-aware attention without position embedding normalization, maintaining separate layer normalization for queries, keys, and values.
Figure 4 illustrates the architecture of our attention blocks. Each block receives the set of tokens $X_{l}$ from the previous layer as input. The first layer receives $X_{1}=T$, which are the tokens produced by patch-size-specific embedding MLPs $E_{r_{0}}..E_{r_{K}}$ that process the foveal patches. Similar to LaPE, we inject relative positional embeddings $\mathrm{PE}$ into every layer.
@@ 3.2 Encoder
Figure 4: Architecture of the Attention block in FLIP. Positional embeddings influence queries, keys and values independent of feature normalization.
The relative positional embeddings $(\hat{x}_{i},\hat{y}_{i}$) are computed with respect to the 2D Gaussian input query. From $\Sigma$, we extract rotation $(\theta_{a},\theta_{b})$ and scale $(\sigma_{x},\sigma_{y})$ via a standard eigenvalue decomposition of the covariance matrix. For each token at position $(x_{i},y_{i})$, we shift coordinates by $(\mu_{x},\mu_{y})$, rotate by $(\theta_{a},\theta_{b})$, and scale by $(\sigma_{x},\sigma_{y})$. We then use separate MLPs to compute positional embeddings for queries, keys and values:
$\mathrm{PE}_{i}^{q,k,v}$ $=\text{MLP}_{i}^{q,k,v}(\hat{x}_{i},\hat{y}_{i}),$
$\mathbf{Q}^{\prime}_{i}$ $=\mathbf{Q}_{i}+\mathrm{PE}_{i}^{q},$
$\mathbf{K}^{\prime}_{i}$ $=\mathbf{K}_{i}+\mathrm{PE}_{i}^{k},$
$\mathbf{V}^{\prime}_{i}$ $=\mathbf{V}_{i}+\mathrm{PE}_{i}^{v},$
where MLP is a two-layer MLP with SiLU activations that maps 2D coordinates to positional embeddings.
Unlike standard transformers, we apply individual layer normalization to queries, keys, and values before projection, which empirically improved performance over a single layer normalization (see Table 2):
$Q$ $=\text{LN}_{Q}(X_{l})W_{Q},$
$K$ $=\text{LN}_{K}(X_{l})W_{K},$
$V$ $=\text{LN}_{V}(X_{l})W_{V},$
where $\text{LN}_{Q}$, $\text{LN}_{K}$, and $\text{LN}_{V}$ are separate LayerNorm modules.
The attention computation proceeds as:
$Q^{\prime}=Q+\mathrm{PE}_{Q},\quad K^{\prime}$ $=K+\mathrm{PE}_{K},\quad V^{\prime}=V+\mathrm{PE}_{V},$
$\text{Attn}(Q^{\prime},K^{\prime},V)$ $=\text{softmax}\left(\frac{Q^{\prime}K^{\prime T}}{\sqrt{d_{k}}}\right)V^{\prime},$
$Y$ $=X_{l}+\alpha\cdot\text{Attn}(Q^{\prime},K^{\prime},V^{\prime})W_{out},$
where $\alpha$ is a learnable scaling parameter initialized to 0.1 for stable training, and $d_{k}$ is the key dimension.
Each attention block is followed by a residual MLP:
$X_{l+1}=Y+\beta\cdot\text{MLP}(\text{LN}(Y)),$
where $\beta$ is another learnable scaling parameter. The MLP uses a bottleneck design with SiLU activations and expansion factor of 4. Depending on the model configuration, we stack 3, 5, 8, or 24 of these blocks to form the complete encoder.
Figure 5: Architecture of the Pixel-Predictor. Encoder features $X_{E}$ are processed through a residual MLP. The attention mechanism computes keys (K) and values (V) enhanced with patch positional embeddings ($\mathrm{PE}_{i}$), while queries (Q) use positional embeddings derived from the target pixels coordinates ($\mathrm{PE}_{t}$). The attended features pass through a residual MLP and a final linear layer to produce mask predictions $M_{t}$ for each query pixel.
3.3 Pixel-Level Mask Prediction
The Pixel-Predictor generates precise segmentation masks by computing attention between encoded patch features and query pixels. Given the encoder output tokens $X_{e}$, we predict mask values for arbitrary pixel coordinates through a sparse attention mechanism.
The predictor consists of preprocessing and postprocessing residual MLP layers surrounding a specialized attention module. We first preprocess the encoder features:
$\hat{X}_{e}=\text{MLP}_{\text{pre}}(X_{e})$
For each target pixel at coordinates $(x_{t},y_{t})$, we compute relative positional embeddings with respect to the input 2D Gaussian by shifting coordinates by $(\mu_{x},\mu_{y})$, rotating by $(\theta_{a},\theta_{b})$, and scaling by $(\sigma_{x},\sigma_{y})$ to get $(\hat{x_{t}},\hat{y_{t}})$:
$\mathrm{PE}_{t}$ $=\text{MLP}_{t}(\hat{x_{t}},\hat{y_{t}}),$
$\mathrm{PE}_{i}$ $=\text{MLP}_{i}(\hat{x}_{i},\hat{y}_{i}),$
where MLP is a two-layer MLP with SiLU activations that maps 2D coordinates to positional embeddings, and $\mathrm{PE}_{i}$ represents the positional embeddings for each encoded and preprocessed input patch ($\hat{X}_{e}$).
In contrast to the encoder, the attention mechanism in the Pixel-Predictor is not residual. We compute embedding-informed keys and values as follows:
$K_{\text{enhanced}}$ $=\text{MLP}_{K}(\text{concat}[\text{LN}_{K}(\hat{X}_{e}),\mathrm{PE}_{i}]),$
$V_{\text{enhanced}}$ $=\text{LN}_{V}(\hat{X}_{e})W_{V}+\text{MLP}_{V}(\mathrm{PE}_{i})$
For each query pixel, we compute:
$Q_{\text{pixel}}$ $=\mathrm{PE}_{t},$
$\text{feat}_{\text{pixel}}$ $=\text{Attention}(Q_{\text{pixel}},K_{\text{enhanced}},V_{\text{enhanced}}),$
$M_{x,y}$ $=\text{Linear}(\text{feat}_{\text{pixel}}+\text{MLP}_{\text{post}}(\text{feat}_{\text{pixel}}))$
where $M_{x,y}$ is the predicted mask logit at position $(x,y)$.